# Unsocial Intelligence: a Pluralistic, Democratic, and Participatory Investigation of AGI Discourse

### v2

@@ Abstract.
Dreams of machines that rival human intelligence have shaped the field of AI since its inception. Yet there remains no agreed-upon conception of what human-level AI or artificial general intelligence (AGI) means. We investigate key social, political, and ethical assumptions made by influential conceptions of AGI and human-level AI. We then draw on feminist, STS, and social science scholarship on the political and social character of intelligence in both humans and machines to defend a pluralistic, democratic, and participatory conception of the topic. We argue that framing AGI or human-level AI as a technical or value-neutral topic leads to political, ethical, and epistemic harm. AGI should not be developed without explicit attention to the values they encode, the people they include or exclude, and a view toward epistemic justice.
@@ 1. Introduction
There is no agreed-upon definition of artificial general intelligence (AGI) (Morris et al., 2023). Yet from influential AI companies (DeepMind, 2022; OpenAI, 2018; Meaker, 2023; Ingram, 2024) and AI researchers (Agüera y Arcas and Norvig, 2023; OpenAI, 2018; Morris et al., 2023; Chollet, 2019; Bubeck et al., 2023) to the increasingly public worry (of Life Institute, 2023; for AI Safety, [n. d.]) about existential risks (Bostrom, 2014; McLean et al., 2023; Naudé and Barten, 2023; Noy and Uher, 2022), AGI and human-level AI have become one of the dominant ways to imagine the transformative potential of AI. At the same time, AGI has had its own skeptics (LeCun, 2022; Toews, 2022; Heaven, 2020). This position paper is shaped by the conviction that what counts as a transformative technology and what makes potentially transformative technologies worth pursuing are fundamentally political and social questions. But current approaches towards AGI risk mistaking these fundamentally social and political questions for technical questions.
@@
The force of some worries about AGI, such as those about existential risks, presumes that there is homogeneity among the views targeted by the worry. We don’t see the lack of homogeneity in current conceptions of AGI as a bug. In our view, it is a crucial feature of the underlying topic: what might it mean for machines to have intelligence that rivals human intelligence? We argue that substantive disagreements about the topic come with the value-laden character of both intelligence and of AI as a technology. By value-laden, we mean that political, social, and ethical values can and should shape conceptions of intelligence, technology, and their intersection. Moreover, without political processes that ensure the legitimacy of those values for those affected by AI, there is likely no way to escape legitimate dissent about what those values should be.
@@
Our paper begins by providing a framework for thinking of AGI as inheriting the value-laden features of both intelligence and technology (Section 2). From there, we move on to investigate the value-laden choices made by influential accounts of AGI and human-level AI (Section 3). We do not frame these choices as inherently misguided. Rather, we see them as occasions to investigate the heterogeneous range of questions currently being asked about human-level machine intelligence. We then propose pathways for more contextual, democratic, participatory, and reflectively value-laden perspectives on what forms of transformative machine intelligence are worth building (Sections 4 and 5).
@@
In investigating discourse on AGI and human-level AI, we follow Morris et al. (2023) in not restricting our discussion to accounts that use the exact phrase “artificial general intelligence”. We agree with them that this topic is part of the long history of thinking of “[a]chieving human-level ‘intelligence” as the “north-star goal” of the AI field (Morris et al., 2023), dating at least as far back as the 1955 Dartmouth AI Conference (McCarthy et al., 1955). We treat discussions of “human-level AI”, “general AI”, or “AI” (such as in the phrase “strong AI” (Searle, 1980)) as in scope for our investigation.
@@ 2. Between human intelligence and technology: AGI’s dual value-laden pedigrees
Building on STS scholarship on the values embedded in technology (Winner, 1980), research communities like FAccT have taken a deep interest in the political, social, and ethical values embedded both in AI tools and in the practices that surround AI (Fishman and Hancox-Li, 2022; Dotan and Milli, 2020; Birhane et al., 2022b; Scheuerman et al., 2021; Hutchinson et al., 2022; Bommasani, 2022; Denton et al., 2020, 2021; Mathur et al., 2022; Shilton, 2018; Broussard et al., 2019; Green, 2021; Blodgett et al., 2020; Viljoen, 2021; Abebe et al., 2020; Birhane and Guest, 2021; Blili-Hamelin, Borhane and Hancox-Li, 2023; Costanza-Chock, 2020). AGI and human-level AI concern not only existing technologies, but also the technologies that many AI builders, researchers, and organizations dream of building in the future. When companies describe their official “long term aim [as] to solve intelligence, developing more general and capable problem-solving systems, known as artificial general intelligence (AGI)” (DeepMind, 2022), or when thought leaders claim that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” (for AI Safety, [n. d.]), they engage in “the process of negotiating between competing perspectives, values, and goals” (Green, 2021).
We interrogate the political, social, and ethical questions undergirding AGI discourse by paying attention to how different accounts occupy the space of competing visions for the future of transformative technologies. We don’t see value-laden assumptions as a flaw. We see them as choices that admit legitimate disagreement through competing values. Our account of alternative paths forward (Section 4) strives to be reflective and explicit about the political and social assumptions of the visions we call for—such as democracy, epistemic justice, contextualism, and participation. Throughout, we also embrace the value of reflectiveness (Boyarskaya et al., 2020; Prunkl et al., 2021) about political, social, and ethical assumptions. We believe that making value assumptions explicit to ourselves and to others often makes for better individual and collective decisions—we might say, for more intelligent “experiments in living” (Mill, 2003; Anderson, 1991).
@@ 2.1. Intelligence is value-laden because it is a thick concept
Intelligence is what philosophers would call a thick evaluative concept: it includes both descriptive and normative elements (Anderson, 2002; Kirchin, 2013, 2017). It contains descriptive elements about what empirical phenomena fall under the concept of intelligent behavior. But it also contains a normative element, because when we evaluate intelligence we are also evaluating the desirability of certain behaviors. Previous research (Anderson, 2002; Alexandrova and Fabian, 2022; Blili-Hamelin, Borhane and Hancox-Li, 2023; Cave, 2020) has argued that when we try to define or measure thick concepts, ethical values are inevitably embedded in the design decisions we make about the boundaries of the concepts and the measurement methodology. Other examples of thick concepts that have value-laden definitions are health and well-being (Alexandrova and Fabian, 2022).
$2.$ g and its critics
The validity of the concept of general intelligence in humans has been extensively criticized. Here we recap a few of the main critiques to provide some background for our later points about how some definitions of AGI make moves that are subject to similar critiques.
@@ 2.2.2. Weaknesses of factor analysis
The use of factor analysis to discover g as a causal factor has been extensively critiqued (Glymour, 1998; Johnson, 2016). These critics argue that factor analysis, on its own, does not rule out alternative causal structures that could explain the same empirical results. Factor analysis can show that a common cause is a possible explanation the observed statistical patterns. But those same patterns could be explained by many other causal hypotheses. Critics argue that common approaches to g fail to rule out alternative causal hypotheses.
We mention this critique of factor analysis not just to situate the epistemic situation of human intelligence metrics, but also as background for our later discussion (Section 3.2.9) of attempts in AI to find the equivalent of g in AI systems. These attempts are subject to the same critique that factor analysis does not rule out alternative causal hypotheses.
One possible response to this critique of factor analysis is to treat g in a deflationary way: to say that it is not explanatory but is merely a shorthand for the correlation between performances on multiple cognitive tests. In this way g could serve as a shorthand for communicating performance on a group of tests, without committing to it being a real causal factor. We mention this attitude towards g because we will see in Section 3.1 that some formulations of AGI declare a similar deflationary approach: claiming that they are defining common standards for AGI primarily in order to have a clear and operationalizable language.
@@ 2.2.3. Circularity
Another critique of psychometrics work purporting to quantify human intelligence is that it cannot be validated without reference to intelligence tests, and that this move is circular (Summerfield, 2023; Boring, 1923; Richardson, 2017; Popper, 2013). As Boring (1923) put it over 100 years ago, “[I]ntelligence is what the tests test”. Boring (1923)’s worry is not simply that intelligence tests need to be appealed to in validating intelligence tests. Rather, it is that the theoretical construct they purport to measure is itself motivated by setting aside the broad range of connotations of intelligence in favor of focusing on whatever can be measured through intelligence tests (Boring, 1923). We see this critique as pointing to how psychometrics research is insufficiently motivated and under-theorized. We see Warne and Burningham (2019) as an example of recent work that doubles down on disconnecting research on g from any attempt at anchoring psychometrics research in a substantive theoretical conception of intelligence. As with the critique of factor analysis (2.2.2), deflationary approaches that avoid ascribing an explanatory role to g may be a potential response to this worry.  [^1] For an overview of recent approaches to motivating psychomterics research, see Deary (2012).
@@ 2.3. Relevance to AGI
The critiques of g and the recognition of intelligence as a thick concept provide the background for why the definition and measurement of human intelligence is a political and social question. For similar reasons, defining and measuring AGI is also a political and social question. AGI is also a thick concept involving both descriptive (what tasks or abilities fall under its definition) and normative (what counts as good machine behavior) criteria. And, as we argue next, current definitions of AGI also embed political and social values in ways that are often not explicitly acknowledged by their authors. In the next section, we identify the design choices that formulations of AGI make, emphasizing their contingency and the types of values that each choice embeds.
@@ 3.1. Purportedly deflationary accounts of AGI
Some definitions of AGI start out with seemingly neutral goals, such as creating a shared language or common standards for measuring AGI, without (at least at first) specifying any normative goals beyond that. This can be viewed as similar to the deflationary view of g described in Sections 2.2.2 and 2.2.3: where g is simply a correlational factor that you can measure from a battery of tests and does not necessarily represent any underlying causal or explanatory factor. One can adopt standards for the sake of being able to consistently compare systems against one another within those standards without assuming that those standards are the only “real” or “objective” standards.
@@
At first, their goal seems to be purely deflationary—they want some shared, operationalizable terminology so that we know when we are referring to similar phenomena. Morris et al. (2023) are inspired by the Levels of Automation framework (International, 2021) used to grade self-driving car technology—a framework that was also allegedly adopted to enhance communication. [^2] However, see Section 3.3 for a discussion of how this framework also drifted from its initial deflationary goals. Their deflationary approach is further underlined by a willingness to change the definition of AGI in the future if we gain the ability to operationalize things that are currently not operationalizable. For example, they currently want to define AGI in terms of “capabilities rather than processes” because we currently have little insight into the underlying mechanisms of many AI systems. They allow that research into mechanistic interpretability might eventually lead to operationalizable definitions involving processes, which then “may be relevant to future definitions of AGI” (Morris et al., 2023).
@@
We’$l$ see in Section 3.3 that these deflationary accounts depart from their initial declared intentions and end up including some value-laden choices. But first, we’$l$ take a tour of some non-deflationary accounts of AGI.
@@ 3.2. Value-laden choices in AGI definitions: a taxonomy
Having outlined deflationary views on AGI, we now identify some value-laden choices that non-deflationary views make in their conception of AGI or human-level AI. We organize these views by the dimensions on which these choices are made (see Appendix A for a summary of these dimensions with examples). We take many of these choices to be better understood as differences in focus and emphasis, rather than as mutually exclusive assumptions about the correct conception of AGI.
@@ 3.2.1. Economic value
OpenAI’s definition is perhaps the most explicit about its values. But we can find other definitions that embed similar values. Suleyman and Bhaskar (2023), for example, define “Artificial Capable Intelligence” by its ability to pass what he calls the “Modern Turing Test”: the capability to turn a starting pot of \$100,000 of capital into \$1,000,000 over several months. Similarly, Nilsson (2005) proposes an “employment test” to replace the Turing test: “To pass the employment test, AI programs must be able to perform the jobs ordinarily performed by humans. Progress toward human-level AI could then be measured by the fraction of these jobs that can be acceptably performed by machines.”
In contrast, Morris et al. (2023) reject definitions based on economic value alone. They think that benchmark tasks should align with what people value beyond economic value, including metrics that are harder to automate or quantify. Here, in tension with their purported deflationary orientation (see Section 3.1), they explicitly argue for having components of AGI that are less operationalizable (because they are harder to quantify), and the reason they are doing so is value-laden: because they think that there are values beyond economic value that are worthwhile to aim for.
@@ 3.2.2. Embodiment
In contrast, criteria like Wozniak’s coffee test require embodied AI systems. This test, first proposed by Apple founder Steve Wozniak (Company, 2010), tasks the machine with going into an “average” American home and making coffee in it. It has since been included in definitions of AGI (Goertzel et al., 2012; Marcus, 2022). AI critics like Weizenbaum (1976) have also cited the field’s relative inattention to embodiment as a major shortcoming.
The choice of whether to include embodiment is value-laden because it incorporates assumptions about which tasks are valuable. Privileging the mental over the physical is a normative choice, and has normative implications in terms of what types of AI systems are likely to be developed, and their impacts on society.

### v3

# Unsocial Intelligence: an Investigation of the Assumptions of AGI Discourse

@@ Abstract.
Dreams of machines rivaling human intelligence have shaped the field of AI since its inception. Yet, the very meaning of human-level AI or artificial general intelligence (AGI) remains elusive and contested. Definitions of AGI embrace a diverse range of incompatible values and assumptions. Contending with the fractured worldviews of AGI discourse is vital for critiques that pursue different values and futures. To that end, we provide a taxonomy of AGI definitions, laying the ground for examining the key social, political, and ethical assumptions they make. We highlight instances in which these definitions frame AGI or human-level AI as a technical topic and expose the value-laden choices being implicitly made. Drawing on feminist, STS, and social science scholarship on the political and social character of intelligence in both humans and machines, we propose contextual, democratic, and participatory paths to imagining future forms of machine intelligence. The development of future forms of AI must involve explicit attention to the values it encodes, the people it includes or excludes, and a commitment to epistemic justice.
@@ 1. Introduction
There is no agreed-upon definition of artificial general intelligence (AGI) (Morris et al., 2023; Mitchell, 2024). Yet from influential AI companies (DeepMind, 2022; OpenAI, 2018; Meaker, 2023; Ingram, 2024) and AI researchers (Agüera y Arcas and Norvig, 2023; OpenAI, 2018; Morris et al., 2023; Chollet, 2019; Bubeck et al., 2023) to the increasingly public worry (Future of Life Institute, 2023; Center for AI Safety, [n. d.]) about existential risks (Bostrom, 2014; McLean et al., 2023; Naudé and Barten, 2023; Noy and Uher, 2022), AGI and human-level AI have become one of the dominant ways to imagine the future potential of AI. At the same time, AGI has had its own skeptics (LeCun, 2022; Toews, 2022; Heaven, 2020; Gebru and Torres, 2024) who risk taking the target of their critiques to be more homogenous and congruent than it actually is.
@@
The lack of homogeneity in current conceptions of AGI is not a bug. It is a feature of the underlying topic: what might it mean for machines to have intelligence that rivals human intelligence? Substantive disagreements come with the value-laden character of both intelligence and of AI as a technology. By value-laden, we mean that political, social, and ethical values can and should shape conceptions of intelligence, technology, and their intersection.(Blili-Hamelin and Hancox-Li, 2023) We lack consensus on the definition of intelligence because of conflicting values. The same applies to the current lack of consensus on AGI.
Ruha Benjamin warns that “a narrow definition of what even counts as technology or intelligence” threatens the ability of communities to imagine worlds worth building.(Benjamin, 2024) What forms of intelligence and technology are worth imagining are political and social questions. However, current approaches to AGI risk mistaking these questions for technical questions. We lay the groundwork for examining these potential mistakes by analyzing and classifying contingent assumptions underlying different definitions of AGI. Contending with the fractured worldviews among conceptions of AGI plants the seeds for resisting harmfully narrow conceptions of machine intelligence.
@@
Our paper proceeds as follows. We begin by providing a framework for thinking of AGI as inheriting the value-laden features of both intelligence and technology (Section  2). We then investigate the value-laden, contingent choices made by influential accounts of AGI and human-level AI (Section  3). Rather than framing these choices as inherently misguided, we see them as opportunities to investigate the limited but heterogeneous range of questions currently being asked about would-be human-level AI. Finally, we sketch pathways for more contextual, democratically legitimate, and participatory perspectives on what forms of machine intelligence are worth imagining (Sections 4 and 5).
@@
Our investigation is not limited to accounts using the exact phrase “artificial general intelligence”. Like Morris et al. 2023, our topic is the long history of treating “[a]chieving human-level “intelligence” as the “north-star goal” of the AI field (Morris et al., 2023), dating at least as far back as the 1955 Dartmouth AI Conference (McCarthy et al., 1955). Discussions of “human-level AI”, “general AI”, or “AI” (such as in the phrase “strong AI” (Searle, 1980)) fall within the scope of our account.
@@ 2. Between human intelligence and technology: AGI’s dual value-laden pedigrees
Building on STS scholarship on the values embedded in technology (Winner, 1980), research communities like FAccT, AIES, and CHI have taken a deep interest in the political, social, and ethical values embedded both in AI tools and in the practices that surround AI (Fishman and Hancox-Li, 2022; Dotan and Milli, 2020; Birhane et al., 2022b; Scheuerman et al., 2021; Hutchinson et al., 2022; Bommasani, 2022; Denton et al., 2020, 2021; Mathur et al., 2022; Shilton, 2018; Broussard et al., 2019; Green, 2021; Blodgett et al., 2020; Viljoen, 2021; Abebe et al., 2020; Birhane and Guest, 2021; Blili-Hamelin and Hancox-Li, 2023; Costanza-Chock, 2020). AGI and human-level AI concern not only existing technologies, but also the technologies that many AI builders, researchers, and organizations dream of building in the future. When companies describe their official “long term aim [as] to solve intelligence, developing more general and capable problem-solving systems, known as artificial general intelligence (AGI)” (DeepMind, 2022), or when thought leaders claim that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” (Center for AI Safety, [n. d.]), they engage in “the process of negotiating between competing perspectives, values, and goals” (Green, 2021).
We interrogate the political, social, and ethical questions undergirding AGI discourse by paying attention to how different accounts embody competing visions for the future of technology. We don’t see value-laden assumptions as a flaw. We see them as choices that admit legitimate disagreement through competing values. Our account of alternative paths forward (Section 4) strives to be reflective and explicit about the political and social assumptions of the visions we call for—such as democracy, epistemic justice, contextualism, and participation. Throughout, we also embrace the value of reflectiveness (Boyarskaya et al., 2020; Prunkl et al., 2021) about political, social, and ethical assumptions. We believe that making value assumptions explicit to ourselves and to others often makes for better individual and collective decisions—we might say, for more intelligent “experiments in living” (Mill, 2003; Anderson, 1991). Our aim in emphasizing reflectiveness is to point to more legitimate methods of determining the values that we want in AI. When large corporations are the only entities championing the narrative that their visions of AGI will transform everyone’s lives, what and whose values ultimately win out becomes a question of power divorced from legitimacy (Burrell and Metcalf, 2024).
@@ 2.1. Intelligence is value-laden because it is a thick concept
Intelligence is what philosophers would call a thick evaluative concept: it includes both descriptive and normative elements (Anderson, 2002; Kirchin, 2013, 2017). It contains descriptive elements about what empirical phenomena fall under the concept of intelligent behavior. It also contains a normative element. Evaluating intelligence inevitably involves assessing the desirability of specific behaviors. Previous research (Anderson, 2002; Alexandrova and Fabian, 2022; Blili-Hamelin and Hancox-Li, 2023; Cave, 2020) has argued that when we try to define or measure thick concepts, we inevitably embed ethical values in the design decisions we make about the boundaries of the concepts and the measurement methodology. Other examples of thick concepts that have value-laden definitions are health and well-being (Alexandrova and Fabian, 2022). The fact that intelligence is value-laden is at the root of debates over its definition and validity.
The validity of the concept of general intelligence in humans has been extensively criticized. One of the principle targets of these critiques is Spearman’s g: a hidden, not directly observable factor expressing “shared variance across a set of intercorrelating cognitive tasks” (Warne and Burningham, 2019). Here we recap a few of the main critiques to provide some background for our later points about how some definitions of AGI make moves that are subject to similar critiques.
@@ 2.1.2. Weaknesses of factor analysis
The use of factor analysis to discover g as a causal factor has been extensively critiqued (Glymour, 1998; Johnson, 2016). These critics argue that factor analysis, on its own, does not rule out alternative causal structures that could explain the same empirical results. Factor analysis can show that a common cause is a possible explanation of observed statistical patterns. But those same patterns could be explained by many other causal hypotheses.
We mention this critique of factor analysis not just to situate the epistemic status of human intelligence metrics, but also as background for our later discussion (Section 3.2.9) of attempts in AI to find the equivalent of g in AI systems. These attempts are subject to the same critique that factor analysis does not rule out alternative causal hypotheses.
One possible response to this critique of factor analysis is to treat g in a deflationary way: to say that it is not explanatory but is merely a shorthand for the correlation between performances on multiple cognitive tests. In this way, g could serve as a shorthand for communicating performance on a group of tests, without committing to it being a real causal factor. This is analogous to formulations of AGI that strive for a deflationary approach, as we describe in Section 3.1.
@@ 2.1.3. Circularity
Another critique of psychometrics work purporting to quantify human intelligence is that it cannot be validated without reference to intelligence tests, making such measurements circular (Summerfield, 2023; Boring, 1923; Richardson, 2017; Popper, 2013). As Boring (1923) put it over 100 years ago, “[I]ntelligence is what the tests test”. Boring (1923)’s worry is not simply that intelligence tests need to be appealed to in validating intelligence tests. Rather, it is that the theoretical construct they purport to measure is itself motivated by setting aside the broad range of connotations of intelligence in favor of focusing on whatever can be measured through intelligence tests (Boring, 1923). This critique points to how psychometrics research is insufficiently motivated and under-theorized. Warne and Burningham (2019) is an example of recent work that doubles down on disconnecting research on g from any attempt at anchoring psychometrics research in a substantive theoretical conception of intelligence. As with the critique of factor analysis (2.1.2), deflationary approaches that avoid ascribing an explanatory role to g may be a potential response to this worry.  [^1] For an overview of recent approaches to motivating psychomterics research, see Deary (2012).
@@ 2.2. Relevance to AGI
The critiques of g and the recognition of intelligence as a thick concept provide the background for why the definition and measurement of human intelligence is a political and social question. For similar reasons, defining and measuring AGI is also a political and social question. AGI is also a thick concept involving both descriptive (what tasks or abilities fall under its definition) and normative (what counts as good machine behavior) criteria. And, as we argue next, current definitions of AGI embed political and social values in ways that are often not explicitly acknowledged by their authors. In the next section, we identify the design choices that formulations of AGI make, emphasizing their contingency and the types of values that each choice embeds.
@@ 3.1. Purportedly deflationary accounts of AGI
Some definitions of AGI start out with seemingly neutral goals, such as creating a shared language or common standards for measuring AGI, without (at least at first) specifying any normative goals beyond that. This can be viewed as similar to the deflationary view of g described in Section 2.1.2: where g is simply a correlational factor that you can measure from a battery of tests and does not necessarily represent any underlying causal or explanatory factor. One can adopt standards for the sake of being able to consistently compare systems against one another within those standards without assuming that those standards are the only “real” or “objective” standards.
@@
At first, their goal seems to be purely deflationary—they want some shared, operationalizable terminology so that we know when we are referring to similar phenomena. Morris et al. (2023) are inspired by the Levels of Automation framework (SAE International, 2021) used to grade self-driving car technology—a framework that was also allegedly adopted to enhance communication. [^2] However, see Section 3.3 for a discussion of how this framework also drifted from its initial deflationary goals. Their deflationary approach is further underlined by a willingness to change the definition of AGI in the future if we gain the ability to operationalize things that are currently not operationalizable. For example, they currently want to define AGI in terms of “capabilities rather than processes” because we currently have little insight into the underlying mechanisms of many AI systems. They allow that research into mechanistic interpretability might eventually lead to operationalizable definitions involving processes, which then “may be relevant to future definitions of AGI” (Morris et al., 2023).
@@
We’ll see in Section 3.3 that these deflationary accounts depart from their initial declared intentions and end up including some value-laden choices. But first, we’ll take a tour of some non-deflationary accounts of AGI.
@@ 3.2. Value-laden choices in AGI definitions: a taxonomy
Having outlined deflationary views on AGI, we now identify some value-laden choices that non-deflationary views make in their conception of AGI or human-level AI. We organize these views by the dimensions on which these choices are made (see Appendix A for a summary of these dimensions with examples). Many of these choices are better understood as differences in focus and emphasis, rather than as mutually exclusive assumptions about the correct conception of AGI.
@@ 3.2.1. Economic value
OpenAI’s definition is perhaps the most explicit about its values. But we can find other definitions that embed similar values. Suleyman and Bhaskar (2023), for example, define “Artificial Capable Intelligence” by its ability to pass what they call the “Modern Turing Test”: the capability to turn a starting pot of \$100,000 of capital into \$1,000,000 over several months. Similarly, Nilsson (2005) proposes an “employment test” to replace the Turing Test: “To pass the employment test, AI programs must be able to perform the jobs ordinarily performed by humans. Progress toward human-level AI could then be measured by the fraction of these jobs that can be acceptably performed by machines.”
In contrast, Morris et al. (2023) reject definitions based on economic value alone. They think that benchmark tasks should align with what people value beyond economic value, including metrics that are harder to automate or quantify. Here, in tension with their purported deflationary orientation (see Section 3.1), they explicitly argue for having components of AGI that are less operationalizable (because they are harder to quantify). The reason they are doing so is value-laden: because they think that there are values beyond economic value that are worthwhile to aim for.
@@ 3.2.2. Embodiment
In contrast, criteria like Wozniak’s coffee test require embodied AI systems. This test, first proposed by Apple founder Steve Wozniak (Fast Company, 2010), tasks the machine with going into an “average” American home and making coffee in it. It has since been included in definitions of AGI (Goertzel et al., 2012; Marcus, 2022). AI critics like Weizenbaum (1976) have also cited the field’s relative inattention to embodiment as a major shortcoming.
The decision to include or exclude embodiment in the definition of AGI is value-laden, as it incorporates assumptions about which tasks are considered valuable. Privileging the mental over the physical is a normative choice. It also carries normative consequences, such as influencing the types of AI systems likely to be developed and their potential impacts on society.